OpenStudy (anonymous):

find the amount of the discount on a $160 bicycle that is on sale for 40% off.

OpenStudy (anonymous):

160 x .4 = 64 160-64= your answer.

OpenStudy (anonymous):

Thanks! do you have time for more? @ReadForLife392

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I gave you a medal :) thanks for helping me. @ReadForLife392

OpenStudy (anonymous):

No problem. : )

OpenStudy (anonymous):

find the unit price for each option below. Round to the nearest cent when necessary. Indicate which option is the better buy. Option 1: 10 candy bars for $6.75 Option 2 12 candy bars for $7.25

OpenStudy (anonymous):

i have a few more, if thats okay with you?

OpenStudy (anonymous):

@ReadForLife392

OpenStudy (anonymous):

Option One: Ok, So you want to divide it. Money always goes on top. 6.75/10 = 0.68 per candy bar Option Two: Same thing. 7.25/12=0.60 per candy bar. Option Two is the better choice. And yes, I can help with more. :)

OpenStudy (anonymous):

Calculate the final price you would pay to buy jacket that originally sells for $65, but is discounted by 40%; you also have to pay 5% sales tax.

OpenStudy (anonymous):

@ReadForLife392 heres the next one :)

OpenStudy (anonymous):

Discount you subtract and tax you add. So, do you remember what I did with the other discount equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, try to solve that real quick then I'll help you with tax.

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

divide 65 by 40? @ReadForLife392

OpenStudy (anonymous):

Ok. So to get the answer you multiply $65 by 40%, but you have to turn the % to a decimal. So to do the move the invisible decimal over two spaces getting you .4. 65 x .4= 26 65-26= discount price

OpenStudy (anonymous):

So the discount price is what?

OpenStudy (anonymous):

39

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i have 3 more

OpenStudy (anonymous):

Natalie picker 135 berries in 15 minutes. If If she continues picking at that rate, how long will it take her to pick 486 berries?

OpenStudy (anonymous):

@ReadForLife392

OpenStudy (anonymous):

No for tax. Tax is the same process as discount, but instead you add. so 39 x .05 = 1.95.... now 39+1.95= your final answer

OpenStudy (anonymous):

First step: 135 divided by 15 equals 9.... So that is 9 berries per minute. So now for 486 berries, multiply 486 x 9= your final answer

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

for the one before, I got 40.95 @ReadForLife392

OpenStudy (anonymous):

Yes, that is right. :)

OpenStudy (anonymous):

okay, and the other one is 4,374?

OpenStudy (anonymous):

@ReadForLife392

OpenStudy (anonymous):

Wait, no... Hold on

OpenStudy (anonymous):

what is it? @ReadForLife392

OpenStudy (anonymous):

Sorry, you divide 486 and 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So then your answer would be?

OpenStudy (anonymous):

54

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

16/50 = x / 156.25

OpenStudy (anonymous):

What is that for?

OpenStudy (anonymous):

Wait, that is the question.

OpenStudy (anonymous):

Ok, for this one you cross multiply. So, 16 x 156.25 = 2500..... now 50x X= 50x So the new equation would be 50x=2500... now divide 50 from 50 and divide it by 2500 which will leave x=your answer

OpenStudy (anonymous):

its a different question:) @ReadForLife392

OpenStudy (anonymous):

Yeah, and I just posted the answer. :)

OpenStudy (anonymous):

Im confused i thought 2,500 was the answer not x @ReadForLife392

OpenStudy (anonymous):

Ok, so 2500 is what you get from multiplying the two number factors. 16/50 x/156.25 So the top number of the firs equation(16) is multiplied by the bottom number of the second equation(156.25) That is were I got 2500, but then you have to include the other two(50 and x) so you also multiply those to get 50x=2500 An equation like this calls for what x equals, so in you have to divide 50 by 50x which would equal 0x, or x. now that you have x=, you need to divide 50 and 2500 to get the answer of 50... so x=50.

OpenStudy (anonymous):

Do you get it?